OpenStudy (anonymous):

I have a question how is my answer wrong I got 2/21 for this equation but I keep getting a warning it is incorrect.

OpenStudy (anonymous):

For what question?

OpenStudy (anonymous):

equation: 7/3 + -5/7 put into simplest form

OpenStudy (aripotta):

7/3 - 5/7 first find the LCD

OpenStudy (anonymous):

21

OpenStudy (anonymous):

I got 2/21 i keep being told it's wrong do you have any suggestions as of why?

OpenStudy (aripotta):

so (7 * 7)/(3 * 7) - (5 * 3)/(7 * 3) = 49/21 - 15/21 = 34/21

OpenStudy (anonymous):

7/3 + -5/7 is the same as 7/3 - 5/7 These two need a common denominator, the lowest is 21 so the new equation is: 49/21 - 15/21 Leave the denominator the same, but suptract the top (49-15=34) The answer is 34/21 (1 13/21)

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

lol I forgot to do the top part heh now I see what went wrong thank you both :)

OpenStudy (aripotta):

no problemo~